# Cryptanalysis on Lightweight Verifiable Homomorphic Encryption

### v1

@@ Abstract
In CCS 2024, Chatel et al. [CKP+24] introduced two lightweight VHE schemes: Replication Encoding (REP) and Polynomial Encoding (PE). A similar approach to REP was used by Albrecht et al. [ADDG24] in Eurocrypt 2024 to develop a Verifiable Oblivious PRF scheme (vADDG). A key approach in these schemes is to embed specific secret information within HE ciphertexts to verify homomorphic evaluations.
This paper presents efficient attacks that exploit the homomorphic properties of encryption schemes. The one strategy is to retrieve the secret information in encrypted state from the input ciphertexts and then leverage it to modify the resulting ciphertext without being detected by the verification algorithm. The other is to exploit the secret embedding structure for modification of the evaluation function $f$ into $f^{\prime}$ which works well on input values for verification purpose.
Our forgery attack on vADDG achieves a success probability of 70.2% under the suggested 80-bit security parameter. Our attack on REP and PE achieves a probability 1 attack with linear time complexity when using fully homomorphic encryption.
@@ 1 Introduction
Verifiable Computation (VC) is a technique that guarantees the correctness of the result of a computation outsourced by a client [GGP10]. This technique allows the client to detect and prevent erroneous or malicious computations by the server. This ensures the integrity of the computations. On the other hand, Homomorphic Encryption (HE) allows computations to be performed on encrypted data without having to decrypt. HE allows the privacy of sensitive information to be maintained while complex calculations are outsourced to the server. Overall, in outsourced computing, VC ensures the integrity of the computational results, while HE guarantees the privacy of the client’s data. Verifiable Homomorphic Encryption (VHE) is the combination of these two technologies.
A naive approach to VHE is the use of the SNARK or other VC techniques for homomorphic computation on encrypted data [ABPS24]. However, this approach results in a highly inefficient solution due to several complex operations required in HE, such as bootstrapping and relinearization. These operations are difficult to integrate seamlessly with existing VC techniques. Alternatively, other lines of research [CKP+24, ADDG24] have focused on the design of efficient VHE schemes that exploit the confidentiality of homomorphic encryption to achieve verifiability at lower computational cost. In this paper, we review these schemes and show that these attempts have been unsuccessful, indicating the need for further improvements.
@@ 1.1 Lightweight Verifiable Homomorphic Encryptions
For the verifiability extension, vADDG uses the following method: If $\mathcal{B}$ wants to evaluate $\mathbf{x}_{i}$ for $1\leq i\leq\alpha$ with verification, $\mathcal{B}$ generates a vector of length $\gamma=\alpha\nu+\beta,$ consisting of randomly permuted $\nu$ copies of each $\mathbf{x}_{i}$ and $\beta$ verification value (i.e. challenge) $\mathbf{x}^{\star}_{k}$, where the verification values are published by $\mathcal{A}$ with its PRF evaluations $\mathbf{z}^{\star}_{k}$ with zero-knowledge proofs. To verify the integrity of the result after outsourcing, $\mathcal{B}$ first recovers the random permutation, then checks whether the evaluation of $\mathbf{x}^{\star}_{k}$’s matches that of $\mathbf{z}^{\star}_{k}$’s. It also checks that the $\nu$ copies of each $\mathbf{z}_{i}$ all have the same value. Then $\mathcal{B}$ accepts the $\mathbf{z}_{i}$’s as honest results. The suggested parameters for achieving 80-bit security in verification are $(\alpha,\beta,\nu)=(105,10,11).$ A notable feature of this (v)ADDG scheme is that the homomorphic computation of this PRF circuit requires only one level of bootstrapping depth by removing the key-switching key.
@@
Chatel et al. have proposed a novel VHE scheme, called Veritas, which supports all operations in BFV and BGV. Compared to existing baseline HE schemes, Veritas introduces an overhead ranging from a factor of 1 to a two-digit factor, depending on the characteristics of the circuit. The scheme’s main idea is to use random values and their precomputed results: For a given circuit, the client, acting as the verifier, holds random verification values and their precomputed results. The client then encrypts a message along with the verification values using specially designed encoders, namely Replication Encoding and Polynomial Encoding. If the homomorphic computation is correct, the decrypted result will contain both the precomputed results and the desired computational result. To verify the integrity of the result, the client checks whether the computed values match the previously known values. If they do, along with additional verifications, the client accepts the result as valid.
@@
PE, the another encoding method in [CKP+24], encodes a message $\mathbf{m}\in\mathbb{Z}_{t}^{N}$ as follows: the client randomly chooses $\alpha\leftarrow\mathbb{Z}_{t}^{\times}$ and a verification value $\mathbf{v}\leftarrow\mathbb{Z}_{t}^{N}$. Next, the client interpolates the message $\mathbf{m}\in\mathbb{Z}_{t}^{N}$ with $\mathbf{v}\in\mathbb{Z}_{t}^{N}$ at $Y=0$ and $Y=\alpha$, respectively. As a result, the client obtains a linear polynomial $\mathbf{m}+\left(\frac{\mathbf{v}-\mathbf{m}}{\alpha}\right)Y\in\mathbb{Z}_{t}^{N}[Y]$. Then, this polynomial is coefficient-wisely encrypted to a ciphertext polynomial $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y\in\mathcal{R}_{q}^{2}[Y]$. The operations in $\mathcal{R}_{q}^{2}[Y]$ are performed as a polynomial in $Y$, where the operations on the coefficients are homomorphic operations. For verification, the client checks whether $\textsf{Dec}(\mathsf{F}(\alpha))=f(\mathbf{v})$, and if they match, the constant term $\textsf{Dec}(\mathsf{F}(0))$ is accepted as the desired computation result $f(\mathbf{m})$.
@@
1.2 Attacks on Lightweight VHE schemes
@@
For the vADDG scheme, the essential step of the forgery attack is to homomorphically recover the positions in the encrypted state where identical values appear. However, since the ADDG scheme utilizes the TFHE scheme which supports only a single level of bootstrapping depth, it is difficult to carry out the forgery attack using the published verification values. To address this, we propose a method for extracting positions by evaluating a characteristic function. Specifically, the adversary can construct a characteristic function that takes each TFHE ciphertext string as input and outputs an encryption of 00 or $1$ and then adds that output to the ciphertext string itself. If this characteristic function outputs 00 at the verification value and outputs $1$ at one of the computation values, the forgery attack succeeds.
For example, if we allow five inputs per one programmable bootstrapping, the adversary can forge the result with a probability of $\left(\frac{31}{32}\right)^{\beta}\approx 70.2\%$ for $\beta=10$. This probability does not change significantly with a decrease in $\alpha$ or an increase in $\nu$, and it satisfies $\lambda$-bit security only when $\beta$ is greater than $O(2^{k}\cdot\lambda)$, where $k$ is the number of inputs allowed in a single bootstrapping.
@@
1.3 Our Methodology: Homomorphic Cryptography
The attacks presented in this paper rely on neither deep mathematics nor sophisticated arguments, but instead maximize the potential power of homomorphic computation to exploit the weakness of the schemes. A possible reason these attacks have not been identified before is that homomorphic computation has not been previously used for cryptanalysis and, as a result, has not been widely explored in this context. It is valuable to examine our methodology in a more abstract way, as it may offer further applications in the design of VHE.
The fundamental approach to achieving verifiability in HE is to introduce an additional secret value alongside the secret key of HE. Both vADDG and REP adopt an index set $S$ as a secret value, which determines the slot or ciphertext position. Similarly, in the case of PE, the element $\alpha$ from $\mathbb{Z}_{t}^{\times}$ serves as the secret value. If an adversary obtains knowledge of these secret values, i.e., $S$ and $\alpha$, a forgery attack becomes trivial by manipulating the secret encryption. Therefore, these schemes conceal the secret values within ciphertexts.
@@ 1.3 Our Methodology: Homomorphic Cryptography
In any case, even without knowing the secret value itself, the adversary can homomorphically implement a cheating circuit by leveraging the functionalities of homomorphic encryption. One way to prevent the proposed attacks is to use somewhat homomorphic encryption with a restricted circuit depth. If the cheating circuit cannot be evaluated within this homomorphic capacity, the attack fails. The vADDG scheme relies on this property; however, our attack circumvents this limitation by constructing a depth-1 cheating circuit.
@@ 1.4 Related Works
There exist two major security concerns for information: integrity and confidentiality. As data utilization expands beyond mere storage to computational applications, the concepts of integrity and privacy have evolved into Verifiable Computation and Homomorphic Encryption, respectively. Various cryptographic primitives have been proposed to achieve VC, such as SNARK [LHW+25], Homomorphic Mac [CF13, GW13], and Homomorphic Signature [CCRS24]. Meanwhile, since Gentry’s breakthrough discovery [Gen09], HE has seen continuous advancements in both functionality and efficiency. Notable schemes such as BGV, BFV, TFHE, CKKS [BGV12, FV12, CGGI20, CKKS17] have emerged, significantly improving the practical applicability of HE.
As HE technology matures and its deployment becomes increasingly widespread, new challenges have surfaced. One major issue is that most research assumes a semi-honest server model, overlooking the necessity for robust integrity guarantees. Combining integrity into HE is not merely about ensuring the correctness of computational results: Due to the inherent malleability of FHE ciphertexts, in the absence of integrity measures, an active adversary can manipulate ciphertexts to undermine confidentiality [LM21, CCP+24], which is known as IND-CPAD attack. Therefore, ensuring the integrity of homomorphic computations is crucial not only as an extension of verifiable computing but also for achieving malicious security in FHE.
However, adding verifiability to Homomorphic Encryption is not a trivial task. Typical integrity mechanisms, such as SNARK, MACs, or digital signatures, conflict with FHE’s requirement for meaningful ciphertext malleability [ABPS24, BCFK21, FGP14, FNP20, GNSV23]. Theoretically, there exist works on maliciously secure FHE that propose new security notions for FHE capable of ensuring integrity [MN24, BSW11]. However, in practice, either no concrete instantiation satisfying these notions is known, or existing approaches are inefficient. Other approaches focus on verifying the integrity of homomorphic computations by leveraging confidentiality, but they tend to be constrained by weak adversary models. [CKP+24] assumes a covert adversary model, and [ADDG24] assumes an adversary capable of performing bootstrapping only once. There are also related works not discussed in this paper; for example, [SXLS24] aimed to ensure integrity in secure matrix multiplication by employing checksums.
2 Preliminary
@@ 2.1 Homomorphic Encryption
Homomorphic Encryption is stated as follows.
@@
A Homomorphic Encryption is a tuple of PPT algorithms
$\Pi=(\mathsf{KeyGen,Enc_{key},Eval_{evk},Dec_{sk}})$
@@ A Homomorphic Encryption is a tuple of PPT algorithms
  • $(\textsf{pk}{},\textsf{sk}{})\leftarrow{\sf KeyGen}(1^{\lambda},f)$: pk includes evaluation keys evk and possibly includes the public encryption key and
  • $\textsf{ct}{}_{x}\leftarrow{\sf Enc_{key}}(x)$ for ${\sf key}=\textsf{sk}{}$ or pk.
  • $\textsf{ct}{}_{y}\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$
@@
Definition 2 (Correcteness).
Let $\Pi$ be a Homomorphic Encryption scheme. $\Pi$ is Correct if the evaluation on the ciphertext decrypts to the correct result.
$\Pr\left[\textsf{Dec}_{\sf sk}(\textsf{ct}{}_{y})=f(x):\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},f))\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{key}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\textsf{Eval}_{\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]=1.$
Examples and some features of the homomorphic encryption are as follows:
2.1.1 TFHE.
TFHE is a TLWE-based HE scheme that is specialized to evaluating boolean circuits[CGGI20], where the (v)ADDG scheme is instantiated with. Its plaintext space is $\mathbb{Z}_{Q}$ and the message space $\mathbb{Z}_{P}$ is encoded into $\mathbb{Z}_{Q}$ for $P<Q$. Thereafter a plaintext is encrypted into the ciphertext space $\mathcal{C}=\mathbb{Z}_{Q}^{n+1}$ where $n$ is determined by the security parameter. For more details, see [Joy22].
2.1.2 BFV.
BFV is an RLWE-based HE scheme that supports the computation of integers [FV12], and it is the scheme in which Veritas is instantiated. Its message space is $\mathbb{Z}_{t}^{N}$, where $t=p^{r}$ is a power of a prime $p$. The ciphertext space is $\mathcal{R}^{2}_{q}=\mathbb{Z}_{q}[X]/(\Phi_{M}(X))$ for $q>t$, where $\Phi_{M}(X)\in\mathbb{Z}[X]$ is the $M$-th cyclotomic polynomial with degree $N=\varphi(M)$. One of the notable features of this scheme is that it supports rotation, $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which is the shifting action on $\mathbb{Z}_{t}^{N}$ with the aid of the rotation key $\textsf{rtk}{}_{i}$. Also, the multiplication between two BFV ciphertexts relies on the relinearization key rlk. For more details, see [GV23].
@@
BFV bootstrapping requires evaluation keys $\textsf{rtk}{}_{1}$ and rlk. Also, TFHE bootstrapping requires a bootstrapping key $\sf{btk}$ for evaluation, which is an encryption of the secret key sk under a different secret key $\textsf{sk}{}^{\prime}$. This bootstrapping key transforms a ciphertext under sk into a ciphertext under $\textsf{sk}{}^{\prime}$, and a ciphertext under $\textsf{sk}{}^{\prime}$ can be switched back to one under sk by using the key-switching key $\sf{ksk}$.
Programmable Bootstrapping.
In TFHE, bootstrapping has an additional feature beyond error refreshing. A lookup table evaluation can be implemented during the bootstrapping process. This bootstrapping technique is called Programmable Bootstrapping (PBS) or functional bootstrapping. PBS enables the homomorphic evaluation of non-linear functions in TFHE.

### v2

@@ Abstract
At CCS, Chatel et al. [CKP+24] introduced two lightweight VHE schemes: Replication Encoding (REP) and Polynomial Encoding (PE). A similar approach to REP was used by Albrecht et al. [ADDG24] in Eurocrypt to develop a Verifiable Oblivious PRF scheme (vADDG). A key approach in these schemes is to embed specific secret information within HE ciphertexts to verify homomorphic evaluations.
This paper presents efficient attacks that exploit the homomorphic properties of encryption schemes. The one strategy is to retrieve the secret information in encrypted state from the input ciphertexts and then leverage it to modify the resulting ciphertext without being detected by the verification algorithm. The other is to exploit the secret embedding structure to modify the evaluation function $f$ into $f^{\prime}$ which works well on input values for verification purposes.
Our forgery attack on vADDG demonstrates that the proposed 80-bit security parameters in fact offer less than 10-bits of concrete security. Our attack on REP and PE achieves a probability 1 attack with linear time complexity when using fully homomorphic encryption.
@@ 1 Introduction
Verifiable Computation (VC) is a technique that guarantees the correctness of the result of a computation outsourced by a client [GGP10]. This technique allows the client to detect and prevent erroneous or malicious computations by the server. This ensures the integrity of the computations. On the other hand, Homomorphic Encryption (HE) allows computations to be performed on encrypted data without having to decrypt. HE allows the privacy of sensitive information to be maintained while complex calculations are outsourced to the server. Overall, in outsourced computing, VC ensures the integrity of the computational results, while HE guarantees the privacy of the client’s data. Verifiable Homomorphic Encryption (VHE) is the combination of these two technologies. A naive approach to VHE is the use of the SNARK or other VC techniques for homomorphic computation on encrypted data [ABPS24]. However, this approach results in a highly inefficient solution due to several complex operations required in HE, such as bootstrapping and relinearization. These operations are difficult to integrate seamlessly with existing VC techniques.
Alternatively, other lines of research [CKP+24, ADDG24] have explored the design of efficient verifiable homomorphic encryption schemes without using general VC techniques. These schemes aim to achieve verifiability at lower computational cost by concealing specific secret informations within the ciphertext and leveraging their confidentiality to ensure integrity. In this paper, we review such ‘lightweight’ verifiable homomorphic encryption schemes and demonstrate that these approaches have been unsuccessful, highlighting the need for further investigation and improvement.
@@ 1.1 Lightweight Verifiable Homomorphic Encryptions
For the verifiability extension, vADDG uses the following method: First $\mathcal{A}$ publishes $\kappa$ verification values (i.e. challenge) $\mathbf{x}^{\star}_{k}$ for $1\leq k\leq\kappa$ and their PRF evaluations $\{\mathbf{z}^{\star}_{k}\}$ with zero-knowledge proofs. If $\mathcal{B}$ wants to evaluate $\mathbf{x}_{i}$ for $1\leq i\leq\alpha$ with verification, $\mathcal{B}$ generates a vector of length $\gamma=\alpha\nu+\beta,$ consisting of randomly permuted $\nu$ copies of each $\mathbf{x}_{i}$ and $\beta$ verification values from $\{\mathbf{x}^{\star}_{k}\}_{1\leq k\leq\kappa}$. To verify the integrity of the result after outsourcing, $\mathcal{B}$ first recovers the random permutation, then checks whether the evaluation of $\mathbf{x}^{\star}_{k}$’s matches that of $\mathbf{z}^{\star}_{k}$’s. It also checks that the $\nu$ copies of each $\mathbf{z}_{i}$ all have the same value. Then $\mathcal{B}$ accepts the $\mathbf{z}_{i}$’s as honest results. The suggested parameters for achieving 80-bit security in verification are $(\alpha,\beta,\nu)=(105,10,11).$ A notable feature of this (v)ADDG scheme is that the homomorphic computation of this PRF circuit requires only one level of bootstrapping depth by removing the key-switching key.
@@
Chatel et al. [CKP+24] proposed a novel VHE scheme, called Veritas, which supports all operations in BFV and BGV. Compared to the baseline HE schemes, Veritas introduces an overhead ranging from a factor of 1 to a two-digit factor, depending on the characteristics of the circuit. The scheme’s main idea is to use random values and their precomputed results: For a given circuit, the client, acting as the verifier, holds random verification values and their precomputed results. The client then encrypts a message along with the verification values using specially designed encoders, namely Replication Encoding (REP) and Polynomial Encoding (PE). If the homomorphic computation is correct, the decrypted result will contain both the precomputed results and the desired computational result. To verify the integrity of the result, the client checks whether the computed values match the previously known values. If they do, along with additional verifications, the client accepts the result as valid.
@@
PE, the other encoding method in [CKP+24], encodes a message $\mathbf{m}\in\mathbb{Z}_{t}^{N}$ as follows: the client randomly chooses $\alpha\leftarrow\mathbb{Z}_{t}^{\times}$ and a verification value $\mathbf{v}\leftarrow\mathbb{Z}_{t}^{N}$. Next, the client interpolates $\mathbf{m},\mathbf{v}$ at $Y=0$ and $Y=\alpha$, respectively. As a result, the client obtains a linear polynomial $\mathbf{m}+\left(\frac{\mathbf{v}-\mathbf{m}}{\alpha}\right)Y\in\mathbb{Z}_{t}^{N}[Y]$. Then, this polynomial is coefficient-wisely encrypted to a ciphertext polynomial $\textsf{ct}{}_{0}+\textsf{ct}{}_{1}Y\in\mathcal{R}_{q}^{2}[Y]$. The operations in $\mathcal{R}_{q}^{2}[Y]$ are performed as a polynomial in $Y$, where the operations on the coefficients are homomorphic operations. For verification, the client checks whether $\textsf{Dec}(\mathsf{F}(\alpha))=f(\mathbf{v})$, and if they match, the constant term $\textsf{Dec}(\mathsf{F}(0))$ is accepted as the desired computation result $f(\mathbf{m})$.
@@
1.2 Attack Description on Lightweight VHE schemes
@@
For the vADDG scheme, the essential step of the forgery attack is to homomorphically recover the positions in the encrypted state where identical values appear. However, since the ADDG scheme utilizes the TFHE scheme which supports only a single level of bootstrapping depth, it is difficult to carry out the forgery attack using the published verification values. To address this, we propose a method for extracting positions by evaluating a characteristic function. Specifically, the adversary can construct a characteristic function that takes each TFHE ciphertext string as input and outputs an encryption of 00 or $1$ and then adds that output to the ciphertext string itself. The forgery attack succeeds if this characteristic function outputs 00 at the verification value and outputs $1$ at one of the computation values.
Using this approach, we can construct a circuit that achieves forgery with probability approximately $2^{-\beta}=2^{-10}$ under the given 80-bit security parameter $(\alpha,\beta,\gamma)=(105,10,11)$. This probability does not change significantly with a decrease in $\alpha$ or an increase in $\nu$, and it satisfies $\lambda$-bit security only when $\beta>\lambda$.
@@
1.3 Methodology: Homomorphic Cryptanalysis
The attacks presented in this paper rely on neither deep mathematics nor sophisticated arguments, but instead maximize the potential power of homomorphic computation to exploit the weakness of the schemes. It is valuable to examine our methodology in a more abstract way, as it may offer further applications in the design of VHE.
The fundamental approach to achieving verifiability in these lightweight schemes is to introduce an additional secret value alongside the secret key of HE. Both vADDG and REP adopt an index set $S$ as a secret value, which determines the slot or ciphertext position. Similarly, in the case of PE, the element $\alpha$ from $\mathbb{Z}_{t}^{\times}$ serves as the secret value. If an adversary obtains knowledge of these secret values, i.e., $S$ and $\alpha$, a forgery attack becomes trivial by manipulating the secret encryption. Therefore, these schemes conceal the secret values within ciphertexts.
@@ 1.3 Methodology: Homomorphic Cryptanalysis
In any case, even without knowing the secret value itself, the adversary can homomorphically implement a cheating circuit by leveraging the functionalities of homomorphic encryption. If the cheating circuit cannot be evaluated within this homomorphic capacity, the attack fails. The vADDG scheme relies on this property; however, our attack circumvents this limitation by constructing a depth-1 cheating circuit.
@@ 1.4 Related Works
There are two major security concerns in information security: integrity and confidentiality. As data utilization expands beyond mere storage to computational applications, these concerns have evolved into Verifiable Computation (VC) and Homomorphic Encryption (HE), respectively. To achieve VC, various cryptographic primitives have been proposed, including SNARKs [LHW+25], Homomorphic MACs [CF13, GW13], and Homomorphic Signatures [CCRS24]. Meanwhile, since Gentry’s groundbreaking discovery [Gen09], HE has seen continuous advancements in both functionality and efficiency. Notable schemes such as BGV, BFV, TFHE, and CKKS [BGV12, FV12, CGGI20, CKKS17] have significantly improved HE’s practical applicability.
A natural follow-up question is whether VC and HE can be integrated, allowing both concerns to be addressed simultaneously. The client would be able to ensure not only the privacy of their data but also the integrity of the computation results. However, integrating verifiability into HE is not merely about verifying the correctness of homomorphic computations. Due to the inherent malleability of FHE ciphertexts, an active adversary—if integrity measures are absent—can manipulate ciphertexts to intentionally trigger incorrect decryption, leading to the IND-CPAD attack and undermining confidentiality [LM21, CCP+24]. Therefore, ensuring the integrity of homomorphic computations is not just an extension of verifiability but a critical requirement for achieving malicious security.
However, adding verifiability to HE is a non-trivial task. It is not even clear what the precise definition of VFHE should be. Although there exists research on maliciously secure FHE that proposes new security notions for ensuring integrity [MN24, BSW11], a well-established and standardized definition has yet to be clearly formulated. In practice, there are two major ways to construct Verifiable HE. One way is to apply integrity mechanisms to the ciphertext operations of HE. However, conventional integrity mechanisms such as SNARKs conflict with HE’s non-arithmetic ciphertext operations. [ABPS24, BCFK21, FGP14, FNP20, GNSV23]. Another research direction attempts to achieve verifiability by generating SNARK-based proof in encrypted computation [ACGaSV24, GGW24, GBK+24]. The state-of-the-art study in this field [GBK+24] claims that its approach is four orders of magnitude faster than methods that generate proof for heavy HE operations. However, these approaches fail to prevent IND-CPAD attacks since integrity can only be verified after decryption.
The verifiable homomorphic encryption scheme targeted in this paper follows a different approach from the two major existing methods. These prior approaches aim to ensure integrity by leveraging the confidentiality of HE and by modifying the message encoding structure of existing HE schemes [CKP+24, ADDG24, SXLS24]. However, there is no well-established hardness assumption or security proof that justifies a reduction from the confidentiality of HE to its integrity, making the connection between the two notions unclear. Thus, these approaches are constrained by weaker adversary models: [CKP+24] assumes a covert adversary model, while [ADDG24] assumes an adversary capable of performing bootstrapping only once.
Our attack leverages the homomorphic properties of HE to forge the underlying encoding structure. This idea was implicitly suggested in [CHLR18], and later formalized in the construction of vADDG [ADDG24], where the authors proposed a naive hardness assumption: that constructing such a cheating circuit within shallow depth is infeasible. This suggests the need for a deeper exploration of what kinds of computations are possible in various homomorphic encryption settings.
2 Preliminaries
@@ 2.1 Homomorphic Encryption
Let $\mathcal{M}$ be a plaintext space and let $\mathcal{F}\subset\{f_{\alpha}\mid f_{\alpha}:\mathcal{M}^{n_{\alpha}}\to\mathcal{M}\}$ be a set of functions on tuples of $\mathcal{M}$. Then Homomorphic Encryption (over $\mathcal{F}$) is stated as follows.
@@
A Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
$\Pi=(\mathsf{KeyGen,Enc,Eval,Dec})$
@@ A Homomorphic Encryption over $\mathcal{F}$ is a tuple of PPT algorithms
  • $(\textsf{pk}{},\textsf{sk}{})\leftarrow{\sf KeyGen}(1^{\lambda},\mathcal{F})$: pk includes evaluation keys evk.
  • $\textsf{ct}{}_{x}\leftarrow{\sf Enc_{pk}}(x)$ for $x\in\mathcal{M}$
  • $\textsf{ct}{}_{y}\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})$ for $y=f(x)$ where $f\in\mathcal{F}.$
@@
Definition 2 ($\delta$-Correcteness).
Let $\Pi$ be a Homomorphic Encryption scheme. $\Pi$ is $\delta$-Correct if the evaluation on the ciphertext decrypts to the correct result.
$\Pr\left[\textsf{Dec}_{\sf sk}(\textsf{ct}{}_{y})=f(x):\begin{aligned} (\textsf{pk}{},\textsf{sk}{})&\leftarrow\mathsf{KeyGen}(1^{\lambda},\mathcal{F})\\ \textsf{ct}{}_{x}&\leftarrow\textsf{Enc}_{\mathsf{\textsf{pk}{}}}(x)\\ \textsf{ct}{}_{y}&\leftarrow\textsf{Eval}_{f,\textsf{evk}}(\textsf{ct}{}_{x})\end{aligned}\right]\geq 1-\delta.$
@@
2.1.2 TFHE.
TFHE [CGGI20] is a FHE scheme that supports gate-level bootstrapping, where the (v)ADDG scheme is instantiated with. Its plaintext space is $\mathbb{Z}_{Q}$ and the message space $\mathbb{Z}_{P}$ is encoded into $\mathbb{Z}_{Q}$ for $P<Q$. Thereafter a plaintext is encrypted into the ciphertext space $\mathcal{C}=\mathbb{Z}_{Q}^{n+1}$ where $n$ is determined by the security parameter.
TFHE bootstrapping requires a bootstrapping key $\sf{btk}$ for evaluation, which is an encryption of the secret key sk under a different secret key $\textsf{sk}{}^{\prime}$. This bootstrapping key transforms a ciphertext under sk into a ciphertext under $\textsf{sk}{}^{\prime}$, and a ciphertext under $\textsf{sk}{}^{\prime}$ can be switched back to one under sk by using the key-switching key $\sf{ksk}$. The TFHE bootstrapping has an additional feature beyond error refreshing. A lookup table evaluation can be implemented during the bootstrapping process. This bootstrapping technique is called Programmable Bootstrapping (PBS) or functional bootstrapping. PBS enables the homomorphic evaluation of non-linear functions in TFHE. For more details, see [Joy22].
2.1.3 BFV.
BFV is an RLWE-based HE scheme that supports the computation of vectors of integers [Bra12, FV12], and it is the scheme in which Veritas is instantiated. Its message space is $\mathbb{Z}_{t}^{N}$, where $t=p^{r}$ is a power of a prime $p$. The ciphertext space is $\mathcal{R}^{2}_{q}=\mathbb{Z}_{q}[X]/(\Phi_{M}(X))$ for $q>t$, where $\Phi_{M}(X)\in\mathbb{Z}[X]$ is the $M$-th cyclotomic polynomial with degree $N=\varphi(M)$. Its multiplication is the Hadamard multiplication. One of the notable features of this scheme is that it supports rotation, $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which is the shifting action on $\mathbb{Z}_{t}^{N}$ with the aid of the rotation key $\textsf{rtk}{}_{i}$. Also, the multiplication between two BFV ciphertexts relies on the relinearization key rlk. As far as known, BFV bootstrapping requires the evaluation keys $\textsf{rtk}{}_{1}$ and rlk.
The BFV scheme is a RLWE based homomorphic encryption scheme that enables computations over vectors of integers. It serves as the underlying encryption scheme for Veritas. The message space of BFV is $\mathbb{Z}_{t}^{N}$, where $t=p^{r}$ denotes a prime power. Its ciphertext space is defined as $\mathcal{R}^{2}_{q}=\mathbb{Z}_{q}[X]/(\Phi_{M}(X))$, where $\Phi_{M}(X)\in\mathbb{Z}[X]$ is the $M$-th cyclotomic polynomial of degree $N=\varphi(M)$, and $q>t$ is a ciphertext modulus.
Multiplication in BFV is performed component-wise using Hadamard multiplication. A notable feature of BFV is its support for rotation operations $\sigma_{i}:\mathbb{Z}_{t}^{N}\to\mathbb{Z}_{t}^{N}$, which cyclically shift the message vector and require a corresponding rotation key $\textsf{rtk}{}_{i}$. Furthermore, ciphertext-ciphertext multiplications are enabled through the use of a relinearization key rlk. As of current knowledge, bootstrapping in BFV necessitates both the rotation keys $\textsf{rtk}{}_{1}$ and the relinearization key rlk. For further technical details, refer to [GV23].